OpenStudy (anonymous):

A triangle has one side that measures 9 units, another side that measures x, and a third side that measures 2 units more than x. The perimeter is 29 units. Which equation would we use to find the value of x?

OpenStudy (anonymous):

9+2+x+x=29 11+2x=29 2x=29-11 2x=18 x=18/2 x=9